# Qianli Wang1,2   Mingyang Wang4,5,6   Nils Feldhus1,2,8   Simon Ostermann2,3,7 Yuan Cao9   Hinrich Schütze4,6   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin  2German Research Center for Artificial Intelligence (DFKI)

### v1

@@ 5 Results
5.1 Automatic Evaluation
5.1.1 Explainability
Counterfactual Examples.
Table 1 exhibits that quantization generally affects LLMs in generating CFEs, leading to a degradation in CFEs quality, particularly in PPL, which indicates CFEs are less fluent and natural. This effect is especially pronounced for smaller LLMs, whereas larger LLMs are less impacted. Notably, in some cases, quantized models perform better or comparably to full-precision models and larger quantized LLMs overall outperform smaller full-precision LLMs.
Natural Language Explanations.
Table 1 illustrates that full-precision LLMs do not consistently outperform their corresponding quantized counterparts. LLMs with lower precision (e.g., bib4) may generate higher-quality explanations than those with higher precision (bib8). Meanwhile, degradation in TIGERScore is most noticeable in quantized models, indicating that errors occur more frequently in the NLEs generated by these models. Notably, for Qwen2.5-14B, a significant performance gap is observed in $F_{1}$ score with AWQ and in BARTScore with GPTQ. Furthermore, we find that larger quantized LLMs do not consistently outperform smaller full-precision LLMs.
Model Method Accuracy % $\uparrow$ # Correct Model Method Accuracy % $\uparrow$ # Correct
Qwen2.5-7B full 63.25 6133 Qwen2.5-32B full 72.45 7025
bib4 60.72 5887 bib4 70.47 6833
bib8 63.01 6109 bib8 72.40 7020
gptq4 60.10 5827 gptq4 71.25 6908
gptq8 63.22 6130 gptq8 72.39 7019
awq 60.60 5876 awq 71.47 6930
Qwen2.5-14B full 73.08 7086 Qwen2.5-72B full 78.26 7588
bib4 70.33 6819 bib4 76.76 7443
bib8 73.06 7084 bib8 78.22 7584
gptq4 25.20 2443 gptq4 76.77 7444
gptq8 73.03 7081 gptq8 78.32 7594
awq 70.61 6846 awq 76.72 7439
@@ 5.1.1 Explainability
Table 2: Knowledge recall accuracy results (%) and number of correct predictions (out of 9696 queries) on the LRE dataset for Qwen2.5-{7B,14B,32B,72B} models across different quantization methods.
(a) Attention: Landmark on continent
(b) FFN: Landmark on continent
Figure 2: Layer-wise neuron contribution score drop across quantization methods for the landmark on continent relation, comparing attention sublayers (left) and feed-forward sublayers (right).
@@
Figure 3: Analysis on the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods. (a) Average contribution scores of top 300 feed-forward neurons across different quantization methods, showing how each method affects neuron activation patterns. (b) Distribution of high-scoring neurons across model layers, showing the number of neurons that exceed the full-precision model’s 300th neuron score threshold in each layer.
(a) Qwen2.5-7B
(b) Qwen2.5-14B
Figure 4: Difference in the entity recall score (EntRec), consistency score (CnstScore) (both defined in Yang et al. (2024)), and accuracy between the full-precision and bib4 quantized models of Qwen2.5-7B and Qwen2.5-14B, evaluated across all layers.
5.1.2 Interpretability
Knowledge Memorization Analysis.
Table 2 shows that quantization introduces varying degrees of accuracy drop depending on both the model size and quantization method. Notably, the accuracy of AWQ, GPTQ4, and bib4 quantized models consistently drop across all model sizes, whereas GPTQ8 and bib8 models retain performance comparable to the original full-precision models. This effect is especially pronounced in smaller models such as Qwen2.5-{7,14}B. Furthermore, the factual recall accuracy degradation is more severe in relations where the original performance has not yet saturated, indicating that such knowledge is more fragile under quantization.
Neuron-level Attribution. Following the neuron-level method of Yu and Ananiadou (2024), we assign each neuron in Qwen2.5-7B a contribution score equal to the increase in log-probability of the correct answer token that the neuron induces. Comparing these scores before and after quantization allows us to quantify how much individual neurons’ influence changes. We analyze four relations that suffer the notable accuracy drop, landmark on continent, person father, person mother, and person sport position. Results are summarized in Figures 2 and 3; additional relations are reported in App. C.5.
Layer-wise Trends. Figure 2 depicts the drop in aggregate contribution scores for attention (left) and feed-forward (right) sub-layers. All quantization methods exhibit a pronounced decline in the final two layers, indicating that most information loss occurs during the network’s decision stages.
Model Method $r_{1}(e_{1})$uparrow $r_{2}(e_{2})$uparrow $r_{2}(r_{1}(e_{1}))$uparrow
@@ 5.1.2 Interpretability
Qwen2.5-32B full 21.61 36.89 18.07
bib4 21.58 36.85 18.09
bib8 21.58 36.85 18.09
gptq4 16.64 34.90 16.73
gptq8 21.18 36.59 17.70
awq 15.11 29.28 14.65
Qwen2.5-72B full 15.71 25.09 9.17
bib4 16.93 25.93 10.13
bib8 20.73 30.84 11.05
gptq4 20.11 34.96 11.14
gptq8 15.74 25.18 9.15
awq 17.08 23.65 9.10

### v2

@@ 5 Results
5.1 Knowledge Memorization Analysis
Table 2 shows that quantization introduces varying degrees of accuracy drop depending on both the model size and quantization method. Notably, the accuracy drop is more obvious in smaller models within the same model family. Additionally, the accuracy of AWQ, GPTQ4, and bib4 quantized models consistently drop across all model sizes, whereas GPTQ8 and bib8 models retain performance comparable to the full-precision models. Furthermore, relations where full-precision performance has not saturated tend to exhibit more severe factual knowledge recall degradation (Table 4), indicating that such knowledge is more fragile under quantization.
Neuron-level Trends. Following the neuron-level method of yu-ananiadou-2024-neuron, we assign each neuron in the model a contribution score equal to the increase in log-probability of the correct answer token that the neuron induces. Comparing these scores before and after quantization allows us to quantify how much individual neurons’ influence changes. We analyze four relations that suffer a notable accuracy drop: landmark on continent, person father, person mother, and person sport position.
We track the top-300 feed-forward neurons in full-precision model, set $\tau$ as their minimum contribution score, and count how many neurons in quantized models exceed $\tau$. As shown in Figure 2 and 2, the per-layer counts decrease after quantization for both models, with the most obvious reductions in the last layers. Complete results across relations are provided in Figures 7 and 8 in App. C.1.
Layer-wise Trends. Figure 2 depicts the drop in aggregate contribution scores for attention (left) and feed-forward (right) sub-layers. All quantization methods exhibit a pronounced decline in the final two layers on Qwen2.5-7B, this trend is consistent on all relations we investigate, as shown in Figure 5 in App C.1. While on Llama3-8B (see Figure 6), the decline occurs in middle-to-late layers, with an increase in the final layers, indicating that information loss patterns vary across model architectures. This divergence may be ascribable to the different ways in which factual knowledge is stored across model families choe2025autoregressivetransformersrememberfacts.
Collectively, both layer-level and neuron-level analyses reveal that quantization primarily affects the network’s last layers. These findings confirm that quantization degrades the decisive information stored in these late layers, accounting for the factual knowledge recall degradation reported in Table 2.333The complete analysis of both models across different relations is provided in Appendix C.1.
(a) Qwen2.5-7B
(b) Qwen2.5-14B
(c) Llama3-8B
Figure 3: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the GPTQ8 quantized and full-precision models, evaluated across all layers.
5.2 Latent Multi-hop Reasoning Analysis
Quantization affects the first-hop Reasoning the most. Table 3 reveals that quantization substantially affects the first hop $r_{1}(e_{1})$, by as much as $30.08\%$, while its impact on the second hop is minimal, with an average of degradation of only $4.25\%$. The $r_{2}(r_{1}(e_{1}))$ deterioration due to quantization is strongly correlated with the ability to correctly predict the bridge entities $r_{1}(e_{1})$, as indicated by a Spearman’s correlation of $0.93$. Nevertheless, the FKR deterioration is not dramatic and considerably acceptable (e.g., Qwen2.5-7B shows a minor average deterioration of 0.77%).
Quantization Effects are not consistent. Figure 3 illustrates that quantization effects on FKR are largely unpredictable. The quantization effect is heterogeneous across layers and variable across models, particularly across architectures, given that facts are stored in various ways across Llama3 and Qwen2.5 models choe2025autoregressivetransformersrememberfacts. Besides, different quantization approaches often affect the model FKR in different manners (Figure 12). Nevertheless, for a given quantization method with different bit width, the layer-wise impact remains broadly similar. Surprisingly, quantization can occasionally even improve FKR (Figure 3). This effect may be attributed to the regularization effect park-etal-2022-robust or quantization-induced noise li2024investigating, which can inadvertently enhance the model’s capability to recall factual knowledge.
bib largely preserve the factual knowledge recall capability. For Qwen2.5-7B, quantization consistently reduce knowledge recall, with the largest degradations under GPTQ and AWQ (Figure 12). By contrast, bib4/8 degrades one-hop reasoning $r_{1}(e_{1})$ and $r_{2}(e_{2})$, as reflected by modest shifts in EntRec and CnstScore relative to the full-precision model, but leaves two-hop reasoning largely unaffected. For Qwen2.5-14B, bib4/8 likewise outperforms other quantization methods, best preserving and occasionally improving FKR (Figure 14). Moreover, despite its greater layer depth than Qwen2.5-7B, Qwen2.5-14B exhibits a similar layer-wise pattern of bib4/8 quantization effects (Figure 12(e), 14(e)), whereas other quantization methods do not. For Llama3-8B, the effect of different quantization methods are alike, i.e., EntRec and CnstScore of quantized model are lower than full-precision model in shallow layers, but becomes much higher in deeper layers (Figure 15). Among the three LLMs evaluated, Llama3-8B is least affected by quantization in terms of FKR.
6 Conclusion
In this work, we examined the impact of quantization on factual knowledge recall. In general, quantization leads to information loss within the model, which in turn degrades factual knowledge recall; this trend becomes more pronounced for smaller models within each model family. Moreover, LLMs quantized with lower bit precision do not invariably perform worse than those with higher bit precision. Quantization can occasionally even improve factual knowledge recall. While quantization effects vary by model and technique, we observed no performance degradation severe enough to compromise its viability as a compression strategy.
Limitations
Our experimental work is confined to English-language datasets. Consequently, the effectiveness of our experiments in other languages may not be comparable and multilingual factual knowledge recall may be simultaneously affected by the degradation of multilingual capabilities due to quantization marchisio-etal-2024-quantization. Extending experiments to the multilingual settings is considered for future work.
We restrict our experiments to Llama3-8B, Qwen2.5-7B and Qwen2.5-14B for computational feasibility: our gradient-based analyses exceed the available GPU memory for larger models. Between the model families (Qwen, Llama, Mistral, DeepSeek), there are lots of architectural equivalences and similarities, e.g., the same attention (grouped-query attention), position embeddings (RoPE), normalization (RMSNorm) or FFN activation (SwiGLU). We argue that, based on our comprehensive experiments, our results are considerably robust and generalizable given the similar architectures compared to other model families.
Moreover, the scope of our experiments is limited to post-training quantization techniques. Investigating the impact of weight-activation quantization, KV cache compression, or quantization-aware training techniques on factual knowledge recall is counted as future work.
Appendix A Dataset Information
LRE
TwoHop-Fact
TwoHop-Fact (Multi-hop Reasoning) One-hop prompts: ($p_{1}$): The mother of Stevie Wonder is Lula. ($p_{2}$): The singer of ‘Superstition’ is Stevie Wonder. Two-hop prompt: ($p$): The mother of the singer of ‘Superstition’ is Answer: Lura
Figure 4: An example from the TwoHop-Fact dataset.
Appendix B Models & Inference Time
Name Citation Size Precision Link
Llama3 llama3modelcard 8B full https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
Llama3 llama3modelcard 8B GPTQ4 https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-GPTQ
Llama3 llama3modelcard 8B AWQ https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-AWQ
Qwen2.5 qwen2024qwen25technicalreport 7B Full https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
Qwen2.5 qwen2024qwen25technicalreport 7B AWQ https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-AWQ
Qwen2.5 qwen2024qwen25technicalreport 7B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int4
Qwen2.5 qwen2024qwen25technicalreport 7B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int8
Qwen2.5 qwen2024qwen25technicalreport 14B Full https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
Qwen2.5 qwen2024qwen25technicalreport 14B AWQ https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-AWQ
Qwen2.5 qwen2024qwen25technicalreport 14B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int4
Qwen2.5 qwen2024qwen25technicalreport 14B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int8
Table 1: Detailed information about used LLMs in our experiments.
Table 1 presents details of the all LLMs used in our experiments (§3.4), including model sizes, quantization approaches and corresponding URLs from the Hugging Face Hub. All models were directly obtained from the Hugging Face repository. All experiments were conducted using A100 or H100 GPUs. Neuron-level and layer-level attribution can be completed within 10 hours, while LMHR experiments take 30 hours averagely.
Appendix C Additional Experiments
C.1 Knowledge Memorization Analysis
Model Method Accuracy % $\uparrow$ # Correct
Qwen2.5-7B full 63.25 6133
bib4 60.72 5887
bib8 63.01 6109
gptq4 60.10 5827
gptq8 63.22 6130
awq 60.60 5876
Qwen2.5-14B full 73.08 7086
bib4 70.33 6819
bib8 73.06 7084
gptq4 25.20 2443
gptq8 73.03 7081
awq 70.61 6846
Llama3-8B full 77.62 7526
bib4 72.19 7000
bib8 76.95 7461
gptq8 71.39 6922
awq 71.83 6965
@@ C.1 Knowledge Memorization Analysis
Table 2: Knowledge recall accuracy results (%) and number of correct predictions (out of 9696 queries) on the LRE dataset for Qwen2.5-{7B,14B} and Llama3-8B models across different quantization methods.
Table 2 illustrates knowledge recall accuracy results on the LRE dataset for Qwen2.5-{7B,14B} and Llama3-8B models across different quantization methods.
Knowledge Recall Accuracy.
In Table 4, and 5, we provide the per-relation factual recall accuracy on Qwen2.5-7B, Qwen2.5-14B, and Llama3-8B models, respectively. As discussion in Section 5.1, we observe that the accuracy drop is more severe in relations where the original performance has not yet saturated.
Neuron Attribution Analysis.
Here we present the neuron-level knowledge attribution results on relations person father, person mother, person sport position, complementing our results in Section 5.1. Analysis on layer-wise neuron contribution scores are shown in Figure 5 and Figure 6, the analysis on top-300 neurons are given in Figure 9 and Figure 10.
(a) Attention: Person father
(b) FFN: Person father
(c) Attention: Person mother
(d) FFN: Person mother
(e) Attention: Person sport position
(f) FFN: Person sport position
Figure 5: Contribution score drop across quantization methods on Qwen2.5-7B for other relationship types, comparing attention sublayers (left) and feed-forward sublayers (right). Rows show different relationships: (a-b) person father, (c-d) person mother, and (e-f) person sport position. The plots reveal how different quantization methods affect knowledge representation across model layers.
(a) Attention: Person father
(b) FFN: Person father
(c) Attention: Person mother
(d) FFN: Person mother
(e) Attention: Person sport position
(f) FFN: Person sport position
Figure 6: Contribution score drop across quantization methods on Llama3-8B for other relationship types, comparing attention sublayers (left) and feed-forward sublayers (right). Rows show different relationships: (a-b) person father, (c-d) person mother, and (e-f) person sport position. The plots reveal how different quantization methods affect knowledge representation across model layers.
@@
Figure 7: Analysis of the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods applied to Qwen2.5-7B. (a) Average contribution scores of top 300 feed-forward neurons across different quantization methods, showing how each method affects neuron activation patterns. (b) Distribution of high-scoring neurons across layers, showing the number of neurons exceeding the full-precision model’s 300th neuron score threshold in each layer.
(a) Contribution score comparison
(b) Layer distribution
Figure 8: Analysis of the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods applied to Llama3-8B. (a) Average contribution scores of top 300 feed-forward neurons across different quantization methods, showing how each method affects neuron activation patterns. (b) Distribution of high-scoring neurons across layers, showing the number of neurons exceeding the full-precision model’s 300th neuron score threshold in each layer.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Analysis of other relationship types under different quantization methods applied to Qwen2.5-7B. Left column: Average contribution scores of top 300 feed-forward neurons across quantization methods. Right column: Distribution of high-scoring neurons across model layers. Each row represents a different relationship: (a-b) person father, (c-d) person mother, and (e-f) person sport position. These visualizations reveal both the magnitude of contribution score changes and their distribution across the model architecture when applying different quantization techniques to various types of factual knowledge.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 10: Analysis of other relationship types under different quantization methods applied to Llama3-8B. Left column: Average contribution scores of top 300 feed-forward neurons across quantization methods. Right column: Distribution of high-scoring neurons across model layers. Each row represents a different relationship: (a-b) person father, (c-d) person mother, and (e-f) person sport position. These visualizations reveal both the magnitude of contribution score changes and their distribution across the model architecture when applying different quantization techniques to various types of factual knowledge.
C.2 Latent Multi-hop Reasoning
Model Method $r_{1}(e_{1})$ $\uparrow$ $r_{2}(e_{2})$ $\uparrow$ $r_{2}(r_{1}(e_{1}))$ $\uparrow$
@@ C.2 Latent Multi-hop Reasoning
Llama3-8B full 7.79 21.39 4.45
bib4 4.94 27.44 3.14
bib8 5.79 18.27 2.90
gptq8 23.62 40.73 20.94
awq 22.35 37.79 19.56